A penguin swimming with the ocean current in search of food takes 1 hour to travel 9 miles. Its return trip upstream to its starting point takes it 3 hours. What is the penguin's speed in still water?
Answer by checkley77(12844)

D=RT
R=D/T
9/1=9 MPH IS THE SPEED WITH THE CURRENT.
R=9/3=3 MPH. IS THE RETURN SPEED.
(9-3)/2=6/2=3 MPH IS THE SPEED OF THE CURRENT.
PROOF:
9-3=6 ANSWER. MPH FOR THE PENGUIN SWIMMING IN STILL WATER.
3+3=6 DITTO.